# Liouville type theorem and Kinetic Formulation for $2\times 2$ Systems of Conservation Laws

### v1

@@ Abstract.
We first study $\mathbf{L}^{\infty}$ vanishing viscosity solutions to $2\times 2$ systems of conservation laws obtained with the compensated compactness method. We show that, if a uniformly convex entropy exists, these solutions satisfy a pair of kinetic equations (nonlocal in velocity), which are then shown to characterize solutions with finite entropy production. Next, we prove a Liouville-type theorem for genuinely nonlinear systems, which is the main result of the paper. This implies in particular that for every finite entropy solution, every point $(t,x)\in\mathbb{R}^{+}\times\mathbb{R}\setminus\boldsymbol{\mathrm{J}}$ is of vanishing mean oscillation, where $\boldsymbol{\mathrm{J}}\subset\mathbb{R}^{+}\times\mathbb{R}$ is a set of Hausdorff dimension at most 1.
@@ 1. Introduction
The existence of entropy solutions is commonly investigated using relaxation techniques, by approximation schemes (such as front tracking or Glimm scheme), or by approximating the equation adding smoothing viscosity terms. Consider the viscous approximations with identity viscosity matrix: it is well known that if the viscous approximations $\boldsymbol{u}^{\varepsilon}$, solving
@@
In this paper, we first study $\mathbf{L}^{\infty}$ vanishing viscosity solutions to $2\times 2$ systems of conservation laws obtained with the compensated compactness method. We show that, if a uniformly convex entropy exists, these solutions satisfy a pair of kinetic equations nonlocal in the kinetic variable (Theorem 1.1), which are shown to characterize solutions with finite entropy production (Proposition 3.5). Next, we prove a Liouville-type theorem for genuinely nonlinear systems (Theorem 1.2), which is the main result of the paper, stating that global isentropic solutions must be constant. This implies in particular that, for every finite entropy solution, there exists a candidate jump set $\boldsymbol{\mathrm{J}}\subset\mathbb{R}^{+}\times\mathbb{R}$ of Hausdorff dimension at most 1 such that every point $(t,x)\in\mathbb{R}^{+}\times\mathbb{R}\setminus\boldsymbol{\mathrm{J}}$ is of vanishing mean oscillation.
@@ 1.2. Contributions of the present paper
Let $\boldsymbol{u}:\Omega\to\mathcal{U}$ be a vanishing viscosity solution to (2.1), assume that (2.1) admits a uniformly convex entropy, and define $\boldsymbol{\chi}_{\boldsymbol{u}},\boldsymbol{\psi}_{\boldsymbol{u}}$ and $\boldsymbol{\upsilon}_{\boldsymbol{u}},\boldsymbol{\varphi}_{\boldsymbol{u}}$ as in (3.9), (3.10). Then there are locally finite measure $\mu_{0},\mu_{1}\in\mathscr{M}(\Omega\times({\underaccent{\bar}{w}},\bar{w}))$ and $\nu_{0},\nu_{1}\in\mathscr{M}(\Omega\times({\underaccent{\bar}{z}},\bar{z}))$ such that
@@
Moreover $\mu_{1}$ and $\nu_{1}$ are positive measures, and for some constant $C>0$, we have
$({\mathtt{p}_{t,x}})_{\sharp}|\mu_{0}|+({\mathtt{p}_{t,x}})_{\sharp}|\nu_{0}|\leq C\,({\mathtt{p}_{t,x}})_{\sharp}\mu_{1}+({\mathtt{p}_{t,x}})_{\sharp}\nu_{1}.$
Here $\mathtt{p}_{t,x}$ denotes the canonical projection on the $t,x$ variables. We recall that given a measurable map $f:X\to Y$ between measure spaces $X,Y$, for any $\mu\in X\times Y$ the pushforward measure $f_{\sharp}\mu\in\mathscr{M}(X)$ is defined by
$f_{\sharp}\mu(A)=\mu(f^{-1}(A))\qquad\forall\;\text{measurable}\;A\subset X.$
@@
    Kinetic formulations that characterize entropy solutions have been obtained for particular systems, see [LPT94b] for the system of isentropic gas dynamics, or [PT00] for a systems in elasticity. A generalization of the kinetic formulation for the system of isentropic gas dynamics with $\gamma=3$ leads to the multibranch solutions introduced by Brenier & Corrias [BC98], which can be viewed as an example of kinetic formulations in the setting of a very specific system of $n$ conservation laws. Equations (1.6), (1.7) (without assumptions on the sign of $\mu_{1},\nu_{1}$) do not characterize entropy solutions, but rather finite entropy solutions (see Proposition 3.5). Since we do not assume any specific structure on the system, the task of characterizing exactly the class of entropy solutions at the kinetic level seems a challenging topic.
    When considering the physical viscosity, as e.g. in [CP10] for the system of gas dynamics, vanishing viscosity solutions might not have a signed dissipation measure for every convex entropy. Therefore they might be in general only finite entropy solutions, satisfying kinetic formulations similar to the one in Theorem 1.1.
@@
This result holds for any weak isentropic (i.e. satisfying (1.8) for every entropy-entropy flux pair) solution, regardless of whether it can be derived through vanishing viscosity approximations, since isentropic solutions automatically satisfy (1.6), (1.7) with $\mu_{i}=\nu_{i}=0$.
A quite standard consequence of Theorem 1.2 is that, if the system admits a uniformly convex entropy, there is a set $\boldsymbol{\mathrm{J}}\subset\mathbb{R}^{+}\times\mathbb{R}$ with Hausdorff dimension at most $1$, such that every point in $\boldsymbol{\mathrm{J}}^{c}$ is a point of vanishing mean oscillation (VMO), see Theorem 5.2. It was known that such a Liouville-type theorem would have implied the VMO property (see, e.g., [DOW03], [CT11]), but a proof of Theorem 1.2 had been missing for some time. The 1-rectifiability of $\boldsymbol{\mathrm{J}}$ remains a challenging problem. Notice that $\boldsymbol{\mathrm{J}}$ can be defined for a vanishing viscosity solution as soon as a uniformly convex entropy exists (since vanishing viscosity solutions are in particular finite entropy solutions, by Theorem 1.1, Proposition 3.5), and in that case it takes the form:
@@ 2. Preliminaries About Conservation Laws
We have the following well known energy bound, see e.g. [Ser00, Section 9.2]. Assuming the existence of a uniformly convex entropy it follows that if $\boldsymbol{u}^{\epsilon}$ is a family of solutions to (2.5) with $\boldsymbol{u}^{\epsilon}(t,x)\in\mathcal{U}$, then for every compact set $K\subset\Omega$ there is a constant $C_{K}$ such that
@@ 3.2. Kinetic Formulation
We can now prove the result of this section, according to which if a function $\boldsymbol{u}:\Omega\to\mathcal{U}$ is a vanishing viscosity solution, then it satisfies a suitable pair of kinetic-type equations. In the following, given a function $\boldsymbol{u}:\Omega\to\mathcal{U}$, we define the bounded function
@@
We prove the Theorem for (1.6), since (1.7) follows from an entirely symmetrical argument.
@@
    $\mu_{1}^{\epsilon}\,\doteq\,(\mathrm{id},w^{\epsilon})_{\sharp}\Big{[}\boldsymbol{\Theta}[{\phi_{1}(\boldsymbol{u}^{\epsilon})}](\boldsymbol{u}^{\epsilon})\big{(}\sqrt{\epsilon}\partial_{x}{\phi_{1}(\boldsymbol{u}^{\epsilon})}\big{)}^{2}\cdot\mathscr{L}^{2}\Big{]}\in\mathscr{M}^{+}_{t,x,\xi}$
@@
    $|\mu_{1}^{\epsilon}|(K)\leq\sup|\boldsymbol{\Theta}|\cdot\sup|\nabla w|^{2}\cdot\big{(}\sqrt{\epsilon}\partial_{x}\boldsymbol{u}^{\epsilon}\big{)}^{2}=\mathcal{O}(1)\cdot C_{K}$ (3.17)
@@
The main results of this paper are based only on the kinetic formulation (1.6), (1.7), which in turn is equivalent to the following notion of finite entropy solution (Proposition 3.5).
Definition 3.3.
Remark 3.4.
For finite entropy solutions it is defined a candidate jump set. In particular we define
where
Here $\bigvee$ denotes the supremum in the sense of measures (see [AFP00, Definition 1.68]) and $\mathcal{E}$ is the set of entropies $\eta:\mathcal{U}\to\mathbb{R}$ (Definition 2.1). We will prove in Section 4 that for every finite entropy solution every point outside $\boldsymbol{\mathrm{J}}$ is of vanishing mean oscillation.
Proposition 3.5.
Let $\boldsymbol{u}\in\mathbf{L}^{\infty}(\Omega,\,\mathcal{U})$. Then $\boldsymbol{u}$ is a finite entropy solution to (2.1) if and only if there are locally finite measures $\mu_{1},\mu_{0}$, $\nu_{1},\nu_{0}$ such that (1.6), (1.7) hold. Moreover, $\boldsymbol{u}$ is an isentropic solution, i.e. for every smooth entropy pair $\eta,q$ it holds $\mu_{\eta}=0$, if and only if
$\partial_{t}\boldsymbol{\chi}_{\boldsymbol{u}}(t,x,\xi)+\partial_{x}\boldsymbol{\psi}_{\boldsymbol{u}}(t,x,\xi)=0\qquad\text{in $\mathscr{D}^{\prime}$}\big{(}\Omega\times\mathbb{R}\big{)}$ (3.20)
$\partial_{t}\boldsymbol{\upsilon}_{\boldsymbol{u}}(t,x,\zeta)+\partial_{x}\boldsymbol{\varphi}_{\boldsymbol{u}}(t,x,\zeta)=0\qquad\text{in $\mathscr{D}^{\prime}$}\big{(}\Omega\times\mathbb{R}\big{)}$ (3.21)
Remark 3.6.
Defining
and with obvious notation also $\widetilde{\boldsymbol{\upsilon}}_{\boldsymbol{u}}$, $\widetilde{\boldsymbol{\varphi}}_{\boldsymbol{u}}$ one can make a symmetric statement: in particular, $\boldsymbol{u}$ is an isentropic solution if and only if (recall (3.22))
Remark 3.7.
At the moment it is an open problem to determine whether an entropy solution (not necessarily coming from vanishing viscosity) is a finite entropy solution. In the presence of a uniformly convex entropy and of a vanishing viscosity solution, of course uniqueness of entropy solutions would solve the problem, by Theorem 1.1 and the above proposition.
@@
where we define the entropy-entropy flux pair associated to $\varrho$
$\eta_{\varrho}(v)\doteq\int_{\underaccent{\bar}{w}}^{\bar{w}}\rho(\xi)\boldsymbol{\chi}[\xi](\boldsymbol{u})\mathop{}\!\mathrm{d}\xi,\qquad\boldsymbol{q}_{\varrho}(v)\doteq\int_{\underaccent{\bar}{w}}^{\bar{w}}\rho(\xi)\boldsymbol{\psi}[\xi](\boldsymbol{u})\mathop{}\!\mathrm{d}\xi.$
$|L_{\varphi}(\varrho)|\leq C_{U,\varphi}\,\|\varrho\|_{\mathcal{C}^{0}}$
Therefore we obtained that the distribution $T$ satisfies the bounds
By a standard application of the Riesz representation theorem we thus obtain the existence of locally finite measures $\mu_{1},\mu_{0}$ such that (1.6) holds. ∎
Remark 3.8.
We could have directly proved that a vanishing viscosity solution is a finite entropy solution and then used Proposition 3.5 to deduce that it satisfies the kinetic formulation of Theorem 1.1. However in this way we would not have obtained the additional properties on the signs and bounds of $\mu_{1},\nu_{1}$ stated in Theorem 1.1, which will be useful for other applications.

### v2

@@ Abstract.
We study $\mathbf{L}^{\infty}$ entropy solutions to $2\times 2$ systems of conservation laws. We show that, if a uniformly convex entropy exists, these solutions satisfy a pair of kinetic equations (nonlocal in velocity), which are then shown to characterize all solutions with finite entropy production. Next, we prove a Liouville-type theorem for genuinely nonlinear systems, which is the main result of the paper. This implies in particular that for every finite entropy solution, every point $(t,x)\in\mathbb{R}^{+}\times\mathbb{R}\setminus\boldsymbol{\mathrm{J}}$ is of vanishing mean oscillation, where $\boldsymbol{\mathrm{J}}\subset\mathbb{R}^{+}\times\mathbb{R}$ is a set of Hausdorff dimension at most 1.
Key words and phrases:
Systems of conservation laws, Liouville type theorem, kinetic formulation
@@ 1. Introduction
The existence of entropy solutions is commonly investigated using relaxation techniques, by approximation schemes (such as front tracking or Glimm scheme), or by approximating the equation adding smoothing viscosity terms. Consider for example the viscous approximations with identity viscosity matrix: it is well known that if the viscous approximations $\boldsymbol{u}^{\varepsilon}$, solving
@@
In this paper, we first study $\mathbf{L}^{\infty}$ entropy solutions to $2\times 2$ systems of conservation laws. We show that, if a uniformly convex entropy exists, entropy solutions are in particular finite entropy solutions, and we show that the latter are characterized by a pair of kinetic equations nonlocal in the kinetic variable (Theorem 1.3). Next, we prove a Liouville-type theorem for genuinely nonlinear systems (Theorem 1.4), which is the main result of the paper, stating that global isentropic solutions must be constant. This implies in particular that, for every finite entropy solution, there exists a candidate jump set $\boldsymbol{\mathrm{J}}\subset\mathbb{R}^{+}\times\mathbb{R}$ of Hausdorff dimension at most 1 such that every point $(t,x)\in\mathbb{R}^{+}\times\mathbb{R}\setminus\boldsymbol{\mathrm{J}}$ is of vanishing mean oscillation.
@@ 1.2. Contributions of the present paper
The main results of this paper are based only on a kinetic formulation, which in turn is equivalent to the following notion of finite entropy solution.
Definition 1.1.
We say that $\boldsymbol{u}$ is isentropic if $\mu_{\eta}=0$ for all entropy-entropy flux pairs $(\eta,q)\in C^{2}$.
It is a simple observation that entropy solutions are in particular finite entropy solutions, if a uniformly convex entropy exists.
Proposition 1.2.
Let $\boldsymbol{u}:\Omega\subset\mathbb{R}^{+}\times\mathbb{R}\to\mathcal{U}$ be an entropy solution, and assume that there exists a uniformly convex entropy $E:\mathcal{U}\to\mathbb{R}$. Then $\boldsymbol{u}$ is also a finite entropy solution.
Proof.
Let $\eta$ be any $C^{2}$ entropy, and let $\kappa>0$ big enough such that $\eta+\kappa E$ is convex. Then we have
$\partial_{t}E(\boldsymbol{u})+\partial_{x}G(\boldsymbol{u})=\mu_{E},$
$\partial_{t}(\eta+\kappa E)(\boldsymbol{u})+\partial_{x}(q+\kappa G)(\boldsymbol{u})=\mu_{\eta+\kappa E}$
where $G$ is the entropy flux of $E$, and $\mu_{E}$, $\mu_{\eta+\kappa E}$ are locally finite negative measures, because $\boldsymbol{u}$ is entropic. Then
$\partial_{t}\eta(\boldsymbol{u})+\partial_{x}q(\boldsymbol{u})=\mu_{\eta+\kappa E}-\kappa\mu_{E}$
which proves the result. ∎
@@
Let $\boldsymbol{u}:\Omega\to\mathcal{U}$ be a finite entropy solution to (2.1), and define $\boldsymbol{\chi}_{\boldsymbol{u}},\boldsymbol{\psi}_{\boldsymbol{u}}$ and $\boldsymbol{\upsilon}_{\boldsymbol{u}},\boldsymbol{\varphi}_{\boldsymbol{u}}$ as in (3.9), (3.10). Then there are locally finite measure $\mu_{0},\mu_{1}\in\mathscr{M}(\Omega\times({\underaccent{\bar}{w}},\bar{w}))$ and $\nu_{0},\nu_{1}\in\mathscr{M}(\Omega\times({\underaccent{\bar}{z}},\bar{z}))$ such that
@@
Moreover, $\boldsymbol{u}$ is an isentropic solution if and only if (1.7), (1.8) hold with $\mu_{i}=\nu_{i}=0$.
@@
    Kinetic formulations that characterize entropy solutions have been obtained for particular systems, see [LPT94b] for the system of isentropic gas dynamics, or [PT00] for a systems in elasticity. A generalization of the kinetic formulation for the system of isentropic gas dynamics with $\gamma=3$ leads to the multibranch solutions introduced by Brenier & Corrias [BC98], which can be viewed as an example of kinetic formulations in the setting of a very specific system of $n$ conservation laws. Equations (1.7), (1.8) (without assumptions on the sign of $\mu_{1},\nu_{1}$) do not characterize entropy solutions, but rather finite entropy solutions (see Theorem 1.3). Since we do not assume any specific structure on the system, the task of characterizing exactly the class of entropy solutions at the kinetic level seems a challenging topic.
    When considering the physical viscosity, as e.g. in [CP10] for the system of gas dynamics, vanishing viscosity solutions might not have a signed dissipation measure for every convex entropy. Therefore they might be a priori only finite entropy solutions, satisfying kinetic formulations similar to the one in Theorem 1.3.
@@
This result holds for any weak isentropic (i.e. satisfying (1.9) for every entropy-entropy flux pair) solution, regardless of whether a uniformly convex entropy exists, since only the kinetic formulation of Theorem 1.3 is used, and isentropic solutions automatically satisfy (1.7), (1.8) with $\mu_{i}=\nu_{i}=0$.
A quite standard consequence of Theorem 1.4 is that, for any finite entropy solution, there is a set $\boldsymbol{\mathrm{J}}\subset\mathbb{R}^{+}\times\mathbb{R}$ with Hausdorff dimension at most $1$, such that every point in $\boldsymbol{\mathrm{J}}^{c}$ is a point of vanishing mean oscillation (VMO), see Theorem 5.2. It was known that such a Liouville-type theorem would have implied the VMO property (see, e.g., [DOW03], [CT11]), but a proof of Theorem 1.4 had been missing for some time. The 1-rectifiability of $\boldsymbol{\mathrm{J}}$ remains a challenging problem. Notice that $\boldsymbol{\mathrm{J}}$ can be defined for a finite entropy solution and it takes the form:
@@ 2. Preliminaries About Conservation Laws
We have the following well known energy bound, see e.g. [Ser00, Section 9.2]. Assuming the existence of a uniformly convex entropy $E:\mathcal{U}\to\mathbb{R}$, it follows that if $\boldsymbol{u}^{\epsilon}$ is a family of solutions to (2.5) with $\boldsymbol{u}^{\epsilon}(t,x)\in\mathcal{U}$, then for every compact set $K\subset\Omega$ there is a constant $C_{K}$ such that
@@
In fact, more precisely for every $M,T>0$, there holds
$\int_{0}^{T}\int_{-M-L(T-t)}^{M+L(T-t)}\big{(}\sqrt{\epsilon}\,\boldsymbol{u}^{\epsilon}_{x}\big{)}^{2}\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}t\leq C\int_{-M-LT}^{M+LT}E(\boldsymbol{u}(0,x))\mathop{}\!\mathrm{d}x$ (2.7)
where $C,L>0$ are positive constants depending only on $f$ and $E$.
@@ 3.2. Kinetic Formulation
We can now prove the result of this section, Theorem 1.3, according to which a function $\boldsymbol{u}:\Omega\to\mathcal{U}$ is a finite entropy solution if and only if it satisfies a suitable pair of kinetic-type equations. In the following, given a function $\boldsymbol{u}:\Omega\to\mathcal{U}$, we define the bounded function
@@
Remark 3.3.
Defining
and with obvious notation also $\widetilde{\boldsymbol{\upsilon}}_{\boldsymbol{u}}$, $\widetilde{\boldsymbol{\varphi}}_{\boldsymbol{u}}$ one can make a symmetric statement to the one in Theorem 1.3: in particular, $\boldsymbol{u}$ is an isentropic solution if and only if (recall (3.11))
Notice that $\widetilde{\boldsymbol{\chi}}_{\boldsymbol{u}},\widetilde{\boldsymbol{\psi}}_{\boldsymbol{u}}$ are now supported on the epigraph $\{\xi\geq w(t,x)\}$ of the first Riemann invariant (recall equation 3.4), while $\widetilde{\boldsymbol{\upsilon}}_{\boldsymbol{u}}$ and $\widetilde{\boldsymbol{\varphi}}_{\boldsymbol{u}}$ are supported on the epigraph $\{\xi\geq z(t,x)\}$ of the second Riemann invariant.
We now prove the Theorem.
@@
where we define the entropy-entropy flux pair associated to $\varrho$
$\eta_{\varrho}(\boldsymbol{u})\doteq\int_{\underaccent{\bar}{w}}^{\bar{w}}\rho(\xi)\boldsymbol{\chi}[\xi](\boldsymbol{u})\mathop{}\!\mathrm{d}\xi,\qquad\boldsymbol{q}_{\varrho}(\boldsymbol{u})\doteq\int_{\underaccent{\bar}{w}}^{\bar{w}}\rho(\xi)\boldsymbol{\psi}[\xi](\boldsymbol{u})\mathop{}\!\mathrm{d}\xi.$
$|L_{\varphi}(\varrho)|\leq C_{U,\varphi}\,\|\varrho\|_{\mathcal{C}^{0}}$
Therefore we obtained that the distribution $T$ satisfies the bounds
By a standard application of the Riesz representation theorem we thus obtain the existence of locally finite measures $\mu_{1},\mu_{0}$ such that (1.7) holds.
Finally, if $\boldsymbol{u}$ is isentropic, then the distribution $T$ defined in the previous step clearly satisfies $T=0$, and this proves the result. ∎
3.3. Vanishing viscosity solutions
Here we prove that vanishing viscosity solutions enjoy the following additional properties, which will be useful for future applications. This also yields a different, more explicit derivation of the kinetic formulation, with a finer characterization of the dissipation measures $\mu_{i},\nu_{i}$.
Proposition 3.4.
If $\boldsymbol{u}:\Omega\to\mathcal{U}$ is a vanishing viscosity solution and a uniformly convex entropy exists, then $\mu_{1}$ and $\nu_{1}$ in (1.7), (1.8) can be taken to be positive measures, and for some constant $C>0$, we have
$({\mathtt{p}_{t,x}})_{\sharp}|\mu_{0}|+({\mathtt{p}_{t,x}})_{\sharp}|\nu_{0}|\leq C\,({\mathtt{p}_{t,x}})_{\sharp}\mu_{1}+({\mathtt{p}_{t,x}})_{\sharp}\nu_{1}.$ (3.16)
Here $\mathtt{p}_{t,x}$ denotes the canonical projection on the $t,x$ variables. We recall that given a measurable map $f:X\to Y$ between measure spaces $X,Y$, for any $\mu\in X$ the pushforward measure $f_{\sharp}\mu\in\mathscr{M}(Y)$ is defined by
$f_{\sharp}\mu(A)=\mu(f^{-1}(A))\qquad\forall\;\text{measurable}\;A\subset Y.$
Proof.
@@ 3.3. Vanishing viscosity solutions
    $\mu_{1}^{\epsilon}\,\doteq\,(\mathrm{id},w^{\epsilon})_{\sharp}\Big{[}\boldsymbol{\Theta}[{\phi_{1}(\boldsymbol{u}^{\epsilon})}](\boldsymbol{u}^{\epsilon})\big{(}\sqrt{\epsilon}\partial_{x}{\phi_{1}(\boldsymbol{u}^{\epsilon})}\big{)}^{2}\cdot\mathscr{L}^{2}\Big{]}\in\mathscr{M}^{+}_{t,x,\xi}$ (3.23)
@@
    $|\mu_{1}^{\epsilon}|(K\times\mathbb{R})\leq\sup|\boldsymbol{\Theta}|\cdot\sup|\nabla w|^{2}\cdot\int_{K}\big{(}\sqrt{\epsilon}\partial_{x}\boldsymbol{u}^{\epsilon}\big{)}^{2}\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}t=\mathcal{O}(1)\cdot C_{K}$ (3.24)
@@
We notice that the energy bound (2.7) translates into precise bounds for the kinetic measures.
Corollary 3.5.
The measures $\mu_{i},\nu_{i}$, $i=0,1$ constructed in the proof of Theorem 1.3 satisfy for all $M,T>0$
$\int_{0}^{T}\int_{-M-L(T-t)}^{M+L(T-t)}\int_{\mathbb{R}}\mathop{}\!\mathrm{d}|\mu_{i}|\leq C\int_{-M-LT}^{M+LT}E(\boldsymbol{u}(0,x))\mathop{}\!\mathrm{d}x$ (3.25)
$\int_{0}^{T}\int_{-M-L(T-t)}^{M+L(T-t)}\int_{\mathbb{R}}\mathop{}\!\mathrm{d}|\nu_{i}|\leq C\int_{-M-LT}^{M+LT}E(\boldsymbol{u}(0,x))\mathop{}\!\mathrm{d}x.$ (3.26)
@@
Recall that $\mu_{1}^{\epsilon}$ is the weak limit of a sequence $\{\mu_{1}^{\epsilon_{k}}\}_{k}$ defined in (3.23). Then using (3.24) with
$K:=\{(t,x)\;|\;x\in(-M-L(T-t),M+L(T-t)),\quad t\in(0,T)\}$
we obtain
$\mu_{1}^{\epsilon}(K\times\mathbb{R})\leq C\int_{K}(\sqrt{\epsilon}\partial_{x}\boldsymbol{u}^{\epsilon})^{2}\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}t\leq\int_{-M-LT}^{M+LT}E(\boldsymbol{u}(0,x))\mathop{}\!\mathrm{d}x$
where in the last inequality we used (2.7). The corresponding inequality for $\nu_{1}$ is proved symmetrically. Finally, the inequalities for $\mu_{0},\nu_{0}$ immediately follow from (3.16). ∎